I'm dying a slow death in statistics. Please help!
In 2004, a small dealership leased 21 Chevrolet Impalas on 2-year leases. When the cars were returned in 2006, the mileage was recorded (see below). Is the dealer's mean significantly greater than the national average of 30,000 miles for 2-year leased vehicles, using the 10 percent level of significance?
Mileage:
40,060 24,960 14,310 17,370 44,740 44,550 20,250
33,380 24,270 41,740 58,630 35,830 25,750 28,910
25,090 43,380 23,940 43,510 53,680 31,810 36,780

Answer by stanbon(75887)

In 2004, a small dealership leased 21 Chevrolet Impalas on 2-year leases. When the cars were returned in 2006, the mileage was recorded (see below). Is the dealer's mean significantly greater than the national average of 30,000 miles for 2-year leased vehicles, using the 10 percent level of significance?
Mileage:
40,060 24,960 14,310 17,370 44,740 44,550 20,250
33,380 24,270 41,740 58,630 35,830 25,750 28,910
25,090 43,380 23,940 43,510 53,680 31,810 36,780
-----------------------
Note: When you post data you should include the sample
mean and the sample standard deviation.
-----------------------
sample mean : xbar = 33949.52
sample standard deviation : s = 11866.17
--------------------------------
Ho: u= 30000
Ha: u > 30000
-----------------
Critical value for right-tail test with df=20 and alpha=10% : t = 1.325
---
test statistic:
t(33949.52) = (33949.52-30000)/[11866.17/sqrt(21)] = 1.5253
p-value = P(t > 1.5253 with df=20) = 0.0714
------------
Conclusion:
Since the p-value is less than 10%, reject Ho.
The dealer's mean is statistically higher than the national average.
-------------------
Cheers,
Stan H.